OpenStudy (anonymous):

How do I solve y-5=-8/7(x-5) I get lost when multiplying -8/7 by 5???

OpenStudy (anonymous):

By 5 or by -5?

OpenStudy (anonymous):

-8/7(x-5) means (-8/7)*(x-5)?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Anyway, the op has gone...

OpenStudy (anonymous):

i am so confused by this problem, Im trying to write the equation in slope intercept form

OpenStudy (anonymous):

"I get lost when multiplying -8/7 by 5???" Still this part?

OpenStudy (anonymous):

right so I get -8/7x but then do i multiply -8/7 (5/1) and get -40/7?

OpenStudy (anonymous):

If you multiply a minus number by a minus number you get a plus, right? So your working is right but the sign is wrong.

OpenStudy (anonymous):

ok by i worked it right, so it would be 40/7

OpenStudy (anonymous):

Yes, why don't u show me what u have so far?

OpenStudy (anonymous):

so I get y-5=-8/7x+40/7

OpenStudy (anonymous):

Yep, now we're in business, just a bit of rearranging and we are done.

OpenStudy (anonymous):

ok so i know i need to add five to both sides what is the best way to do that?

OpenStudy (anonymous):

Can I just ask, did the question say "solve" or just "simplify" or something else? Because you can't actually solve this equation only arrange it in some form or other.

OpenStudy (anonymous):

it says to write the equation in slope intercept form y=mx+b and then write in in Ax+By=C where A B Care integers and A>0

OpenStudy (anonymous):

Right, back to your adding 5. After we have done that we get y on the left side, which is what we want, right? And on the right side we have -8/7x+40/7 +5. So we have y = -8/7x+40/7 +5 and what I would like you to do now is multiply everything (both sides) by 7.

OpenStudy (anonymous):

7y=-8x+75?

OpenStudy (anonymous):

thank you so much for the help, I figured it out thanks to you!!!

OpenStudy (anonymous):

Anytime..

OpenStudy (anonymous):

how do i change the question status to answered?